Suppose that autonomous consumption is 500, government purchases are 1.500, planned investment spending is 500, net exports are 0, and the MPC is 0.75.


Equilibrium GOP is equal to $_______   
Answer #1

Equilibrium is at

Y=C+I+G+NX

C=consumption =autonomous consumption +MPC*Y

=500+0.75Y

Y=GDP

I=500

G=1500

NX=0=net export

Y=500+0.75Y+500+1500+0

0.25Y=2500

Y=10000

The equilibrium GDP is $10000